# Tabby: Tabular Data Synthesis with Language Models

### v1

@@ E.2 Applying Tabby to Transformer MLPs or Attention Blocks
We examine in-detail the performance of Tabby models with MoE applied to the transformer MLPs or attention blocks. We use the following terminology to refer to these architectures, which are visualized in Figure 8:
@@
We focus on Tabby MH in Sections 3.1-3.4 because it demonstrates top performance in most settings. We display results for the MMLP and MMLP+MH architectures across all six datasets for MLE, discrimination and DCR in Tables 8, 9 and 10, respectively. All three metrics are displayed for the MA architecture on two datasets in Table 11.
@@
| Original (Upper Bound) | ||||||
@@ E.3 Additional Metrics for Experiment Applying Tabby MH to Models of Varying Sizes
The results in Section 3.2 compare the MLE scores of Plain-trained models of varying sizes on the Travel dataset. Table 12 incorporates the results for the Diabetes and House datasets as well. Similarly, Table 13 presents results for models trained using GReaT and Tabula (TapTap is not included here, because TapTap-pretrained checkpoints are available only for Distill-DGPT2 and GPT2).
@@
| Original (Upper Bound) | ||||||
@@
| Original (Upper Bound) | ||||||
@@ E.4 Analysis from Tracking the Adaptation to Individual Columns
We observe that Occupancy is the largest contributor to the model’s loss until step . While Median Income’s loss is initially the second-lowest, it improves little throughout the training process and exhibits the highest loss of all columns at the end of training. Additionally, we view that convergence occurs around step .
@@
E.5 Hyperparameters for All Experiments
We list the learning rates chosen for Section 3.1 in Table 14, Section 3.2 in Table 15 and Section 17 in Table 17. We select the learning rate that yields lowest training loss on one run from the set . For non-LLM methods in our experiments, we use the hyperparameters recommended by their respective papers.
| Diabetes | Travel | Adult | Abalone | Rainfall | House | |
| Plain Non-Tabby | ||||||
| Plain Tabby MMLP | ||||||
| Plain Tabby MH | ||||||
| Plain Tabby MMLP+MH | ||||||
| GReaT Non-Tabby | ||||||
| GReaT Tabby MMLP | ||||||
| GReaT Tabby MH | ||||||
| GReaT Tabby MMLP+MH | ||||||
| GTT Non-Tabby | ||||||
| GTT Tabby MMLP | ||||||
| GTT Tabby MH | ||||||

### v2

@@ E.2 Applying Tabby to Transformer MLPs or Attention Blocks
We examine in-detail the performance of Tabby models with MoE applied to the transformer MLPs or attention blocks. We use the following terminology to refer to these architectures, visualized in Figure 6:
@@
We focus on Tabby MH in Sections 3.1-3.4 because it demonstrates top performance in most settings. We display results for the MMLP and MMLP+MH architectures across six datasets for MLE, discrimination and DCR in Tables 14, 15 and 16, respectively. All three metrics are displayed for the MA architecture on two datasets in Table 17.
@@
| Non-Synthetic (Upper Bound) | ||||||
@@ E.3 Additional Metrics for Experiment Applying Tabby MH to Models of Varying Sizes
The results in Section 3.2 compare the MLE scores of Plain-trained models of varying sizes on the Travel dataset. Table 18 incorporates the results for the Diabetes and House datasets as well. Similarly, Table 19 presents results for models trained using GReaT and Tabula (TapTap is not included here, because TapTap-pretrained checkpoints are available only for Distill-GPT2 and GPT2).
@@
| Non-Synthetic (Upper Bound) | ||||||
@@
| Non-Synthetic (Upper Bound) | ||||||
@@ E.4 Analysis from Tracking the Adaptation to Individual Columns
Individual column losses are shown in Figure 8. We observe that Occupancy is the largest contributor to the model’s loss until step . While Median Income’s loss is initially the second-lowest, it improves little throughout the training process and exhibits the highest loss of all columns at the end of training. Additionally, convergence occurs across most columns around step .
@@
E.5 Investigating Tabby’s Generalization to Unseen and High-Cardinality Categorical Features
We now seek to demonstrate our fifth claim.
Claim 5: Tabby’s language modeling capabilities enable it to capture the underlying semantic structure of column values unseen during pretraining, allowing it to generate novel yet realistic values beyond the pretraining distribution.
Setup: We train a Llama 3-8B Tabby MH model for one epoch (LR ) on a 5000-row synthetic dataset, designed to imitate a common business scenario such as a product database. The dataset includes two columns: “Code” contains a 6-digit random number, prepended by “upi_” or “sku_” and followed by a year [2000,2026). “ID” contains three numbers, where the first number is [1,999] and other numbers are [0,999], followed by a hyphen and a single lowercase or uppercase letter. We ensure that each trainset Code and each ID are unique. Some examples are in Table 20.
| Code | ID |
| upi_400727_2001 | 427.424.604-d |
| sku_305434_2021 | 60.272.433-i |
| sku_871297_2012 | 727.884.985-G |
| upi_712558_2003 | 160.401.237-E |
| sku_911473_2008 | 856.31.239-c |
Results: We find that the model produces examples of the same structure as the training data. In particular, out of 1000 generated samples:
1000/1000 rows meet all of the rules for the Code column value,
999/1000 rows meet all of the rules for the ID column value (one row omitted the final hyphen and letter),
Only one Code value in the trainset reappears in the synthetic set,
Only one synthetic Code value is reused across the synthetic set,
All synthetic IDs are unique and
No trainset IDs appear in the synthetic set.
Discussion: Tabby’s generalization capabilities are particularly valuable in real-world settings where datasets include non-categorical string features such as names, product IDs, addresses, or telephone numbers. While these capabilities extend Tabby’s applicability across diverse data domains, the model’s behavior in such cases depends on the implementation details of its underlying language model and tokenizer.
For example, when out-of-vocabulary (OOV) tokens appear, their handling is determined by the specific tokenizer used. In the Distilled-GPT2 backbone employed in our main experiments, the tokenizer applies a byte-level variant of Byte Pair Encoding (BPE) [gpt2], following the subword tokenization approach of [sennrich2016neural]. This ensures that every possible UTF-8 string can be represented without resorting to dedicated OOV tokens. Instead, previously unseen strings are decomposed into byte-level subtokens, which the model can interpret and adapt to based on surrounding context. By contrast, other tokenizers may rely on explicit OOV symbols or restricted vocabularies, potentially affecting Tabby’s ability to generalize to novel strings and impacting overall performance in such settings.
E.6 Hyperparameters for All Experiments
We list the learning rates chosen for Section 3.1 in Table 21, Section 3.2 in Table 22 and Section 24 in Table 24. We select the learning rate that yields lowest training loss from the set . For non-LLM methods in our experiments, we use the hyperparameters recommended by their respective papers.
| Diabetes | Travel | Adult | Magic | Shoppers | Abalone | Rainfall | House | |
| Plain Non-Tabby | ||||||||
| Plain Tabby MMLP | ||||||||
| Plain Tabby MH | ||||||||
| Plain Tabby MMLP+MH | ||||||||
| GReaT Non-Tabby | ||||||||
| GReaT Tabby MMLP | ||||||||
| GReaT Tabby MH | ||||||||
| GReaT Tabby MMLP+MH | ||||||||
| GTT Non-Tabby | ||||||||
| GTT Tabby MMLP | ||||||||
| GTT Tabby MH | ||||||||